OpenStudy (anonymous):

Choose the equation that could be used to find three consecutive integers whose sum is 36. n + (n + 2) + (n + 4) = 36 n + (n + 1) + (n + 3) = 36 n + (n + 1) + (n + 2) = 36 n + (n − 1) + (n − 3) = 36

OpenStudy (anonymous):

n+(n+1)+(n+2)=36 :o)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

To solve it: n+(n+1)+(n+2)=36 3n+3=36 3n=33 n=11 11+(11+1)+(11+2)=36 11+12+13=36 36=36 Does that make sense? :o)